OpenStudy (anonymous):

PLEASE HELP!!! What is the simplified form of x^2-1/9x/x^2+2x+1/3x^2 a) x(x+1)/3(x-1) b) 3x(x+1)/(x-1) c) x(x-1)/(x-1) d) (x-1)/3x(x+1)

OpenStudy (anonymous):

hold on ill do the equation

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

is that 1/(9x) or x/9, and in the denominator 1/(3x^2) or x^2/3?

OpenStudy (anonymous):

\[\frac{ x ^{2}-1 }{ 9x } \over \frac{ x ^{2}+2x+1 }{ 3x ^{2} } \]

OpenStudy (anonymous):

You should try, http://www.wolframalpha.com/, enter your text into the box

OpenStudy (anonymous):

I got $$\frac{x-1}{3(x+1)}$$

OpenStudy (anonymous):

a) x(x+1)/3(x-1) b) 3x(x+1)/(x-1) c) x(x-1)/3(x+1) d) (x-1)/3x(x+1)

OpenStudy (anonymous):

your answer is closest to d.

OpenStudy (anonymous):

might have made a typo, just type it in yourself on wolframalpha

OpenStudy (anonymous):

i did none of the answers pop up

OpenStudy (anonymous):

scroll down to alternate forms, all of those are equivilent. look for the one factored